OpenStudy (anonymous):

Urgent help! Tomorrow is my exam! Form the differential equation of the family of parabolas having vertex at the origin and axis along positive y-axis

OpenStudy (anonymous):

@No.name @ganeshie8

OpenStudy (anonymous):

 So I know the formula is x^2 = 4ay

ganeshie8 (ganeshie8):

thats a good start !

OpenStudy (anonymous):

So If I do y' I get 2x dy/dx= 4a What do I do after this?

ganeshie8 (ganeshie8):

x^2 = 4ay differentiating both sides with respect to x gives you 2x= 4a*dy/dx

ganeshie8 (ganeshie8):

solve the value of a from first equation and plugin above^^

ganeshie8 (ganeshie8):

x^2 = 4ay a = x^2/4y

ganeshie8 (ganeshie8):

so the differential equation becomes : 2x= 4(x^2/4y)*dy/dx simplify

OpenStudy (anonymous):

I'm a bit confused and very nervous, I should probably avoid learning new concepts a few hours before exam.. right?